Shania bought a $1,550 drum set on an installment plan. The installment agreement included a 20% down payment and 12 monthly payments of $110.78 each. how much is the down payment
1
Expert's answer
2020-12-14T19:44:45-0500

payment for 12 months = "110.78\times 12=\$1329.36"

down payment "=1550-1329.36=\$220.64"